A serving of food contains 12 grams of carbohydrate, 6 grams of protein and 6 grams of fat. Based on this information, a serving of this food supplies

a.) 116 kcal

b.) 120 kcal

c.) 126 kcal

d.) 132 kcal
Answer #1

The correct answer is c.) 126 kcal

Amount of carbohydrates is the food is 12 grams.

energy in 1 gram carbohydrates is = 4 kilocalories

energy present in 12 grams carbohydrates is 12 * 4 = 48 kilocalories

Amount is proteins in the food is 6 grams.

energy in 1 gram protein is = 4 kilocalories

energy present in 6 grams proteins is 6 * 4 = 24 kilocalories

Amount or fats in the food is 6 grams.

energy in 1 gram fat is = 9 kilocalories

energy present in 6 grams fat is 9 * 6 = 54 kilocalories.

Total energy in the serving of the food is = energy from carbohydrates + energy from proteins + energy from fats

which is 48 + 24 + 54 = 126 kilocalories.

This is a brief explanation. Hope this helps you out. Please feel free to give a like and contact us if you have any queries. Thank you have a great day.